OpenStudy (anonymous):

Solve by using the quadratic formula to solve x2 + 6x = –3. a.) -6±√-6 over 6 b.) -3±2√3 c.) -3±√6 d.) -3±√-3

OpenStudy (anonymous):

move the -3 from the right side by adding 3 to both sides of the equation. identify your coefficients and plug into the quadratic formula.

jhonyy9 (jhonyy9):

b. is right

OpenStudy (anonymous):

\[\LARGE x_{1/2}={-b\pm\sqrt{b^2-4ac}\over 2a}\] \[\LARGE ax^2+bx+c=0\]

OpenStudy (anonymous):

I solved to -3±2√3. Thank you.